# MARAGS: A Multi-Adapter System for Multi-Task Retrieval Augmented Generation Question Answering

### v1

@@ 1. Introduction
Retrieval augmented generation (RAG) has been a popular approach for question answering systems for some time (Lewis et al., 2020), although recently has become a very popular approach for a wide range of tasks due to the zero-shot capabilities of large language models (LLMs) with an appropriate prompt and access to the relevant context for the task. Despite the existence of numerous question answering benchmarks, many do accurately reflect the diverse usage of current RAG systems. Thus, tracking both the efficacy of certain RAG architectures as well as tracking process remains difficult. The CRAG (Yang et al., 2024) benchmark aims to resolve this with 3 different subtasks representing realistic RAG usage scenarios. The final key element of the CRAG benchmark is it’s scoring metric which explicitly punishes hallucinations. With the rising capabilities of LLMs, increasingly their outputs are taken at face value, despite the known issue of hallucinations. This has lead to high profile incidents causing concern with their use (law, [n. d.]). The CRAG score aims to punish hallucinated answers and encourages returning missing answers, equivalent to returning ”i don’t know” from the model, by giving scores of 1, 0, and -1 to correct, missing, and hallucinated answers respectively. To address these various tasks, we train individual adapters for each of the various tasks, as well as API call generation required for accessing information in the mock API. This approach allows us to use a single LLama 3 (Dubey et al., 2024) in memory while swapping out adapters based on the current needs.
@@ 2. Related Works
The initial approach of Lewis et al. (Lewis et al., 2020) showed the benefits of providing additional text context for seq2seq models for NLP tasks that are knowledge entensive. Using BART, they were able to improve question answering tasks using dual biencoders for retrieval and training the model jointly, without the need for knowing which documents were relevant.
Adapters have become increasingly used since introduced by Houlsby et al. (Houlsby et al., 2019). LoRa (Hu et al., 2022) has become a popular adapter approach, particularly for LLMs as they have grown substantially larger in recent years. The use of adapters allows modifying a models output without training the entire network, which substantially saves on VRAM memory when training. Hu et al. (Hu et al., 2022) discovered that when replacing a dot product between large vectors with an intermediate dot product of a much lower rank vector, the impacts on performance were minimal while further reducing the training parameters required.
@@ 3. CRAG Dataset
CRAG is question answering dataset aimed at providing a realistic task to benchmark RAG systems as they are used in practice with a diverse set of questions, including 8 distinct question types and 5 distinct domains. Additionally, two sources of diversity which pose difficulty for LLMs are how dynamic a question’s answer is and the popularity of the topic of the question. As shown in the baseline systems, real-time answers pose a challenge for RAG systems and they similarly struggle when the topic of the question is less common (referred to as ”torso” and ”tail” questions).
@@ 3.3. Task 3
Finally, Task 3 represents an extension of Task 2, where the system has access to both HTML and the knowledge graph API. However, in this task, the number of HTML documents that are to be processed is 50. This task is meant to measure both the computational efficiency of the approach as well as its ability to filter large amounts of potentially irrelevant information.
@@ 4.1. Webpage Processing
In our case, we utilize the structure of the HTML to provide the segmentation. We traverse the tree structure of the parse HTML with breath first search. Any time the length of the text contained within a node (which includes all of the text of its decedents) is less than 2000 characters, we treat that as a segment. If a leaf node is reached with greater than 2000 characters, the text is split on the white space and into as many segments are needed such that each segment is under the threshold. The segment length was determined via inspection of of HTML documents and their associated segmentation, thus future work could treat this as a hyperparameter and tune for performance.
@@ 4.2. API Call Generation
We acknowledge that this approach to annotation is not optimal, as it likely results in successful, but incorrectly selected API endpoint calls. However, manually annotating each question to determine the correct API call and validating the returned information were indeed relevant would have been too time consuming give the size of the dataset.
@@ 4.3. Candidate Ranking
For candidate ranking, we attempted 4 different candidate ranking approaches. We utilized TF-IDF, a biencoder, cross-encoder, and an ensemble of mentioned approaches (using mean rank as the ranking metric). Our TF-IDF implementation is based on Scikit-Learn (Pedregosa et al., 2011). The biencoder and cross-encoder are from the SentenceTransformer (Reimers and Gurevych, 2019) library, specifically the ”multi-qa-MiniLM-L6-cos-v1” 111https://huggingface.co/sentence-transformers/multi-qa-MiniLM-L6-cos-v1 and ”ms-marco-MiniLM-L-6-v2” 222https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2 respectively.
Evaluating candidate ranking in isolation is difficult, as relevant information is not labeled, so using the system accuracy and CRAG score is the most straight forward way to compare differences in each system. However, to test the various systems, we use the base Llama-3 8B model with not adapters for each retrieval approach and use the accuracy metric to determine the best performing approach. We use accuracy instead of CRAG at this stage for ranking, as we think this is a better representation of how often the relevant information retrieved. For a test set, we randomly select 500 samples from the Task 1 dataset.
@@
The results of this experiment are shown in Table 1. From the results, the cross-encoder is the best performing system, thus we used it for our retriever. We suspect that with proper tuning TF-IDF would and ensembling would be much more performant overall, but as mentioned running extensive experimentation is difficult as it required LLM generation to get an overall accuracy score. Using an LLM to label passages as relevant or not is a possible approach to allow for tuning of just the retriever, however we did not explore this.
Despite the cross-encoder being the most computationally expensive approach, we found it to be fast enough processing the candidates in the required 30 seconds per sample. In the case of Task 3, we found it necessary to use Python’s multiprocessing library to process multiple HTMLs simultaneously to meet the runtime requirement.
@@ 4.4. Retrieval Augmented Generation
Our initial prompt structure started with the question first, then all of the retrieved candidates prior to the Llama model response, however we noticed that often due to how much context would be provided, the model would occasionally forget the question being asked. For example a question like ”What is the elevation of Honolulu?” would result in an answer of ”Honolulu population is 343,421”, indicating the model remember the subject of the question, but incorrectly answered with the population, rather than elevation. Our subsequent prompt structure placed the question after the context, which resolved the issue.
For training Llama 3, we trained LoRa models for each task individually. Given the penalization of hallucinations in the scoring metric, we try to take steps to mitigate further hallucinations due to finetuning, as it has been observed that finetuning LLMs can be a source of further hallucinations (Gekhman et al., 2024). This likely applies to RAG systems in cases where the expected answer is not answerable given the context provided, i.e. no relevant information is given and the question is not answerable without further context, yet the model is trained to output the target answer regardless. Thus for our training setup, we first relabel the target for training samples in cases where our candidate retrieval system likely does not provide the correct information and Llama does not know the answer without that information. We use the provided dev set for training, with the 500 set sample used for retrieval comparison treated as our holdout set.
@@
- LoRa | 0.398 | 0.602 | -0.204 |
- LoRa (relabeled) | 0.242 | 0.056 | 0.186 |
Our initial approach for determining which training samples need relabeling was has been explored previously (Wang et al., 2018a). A common and simple approach to filter/relabel incorrectly333”Incorrectly” here simply means in the context of our retrieval system, not that the provided answer is not true. labeled samples is to use a particular sample’s training loss after training to the point of overfitting. High loss examples after overfitting likely indicate examples that are incorrectly labeled and thus can be filtered out. Not all samples with high loss will be incorrect labels, instead simply being hard examples, yet typically the benefit of disposing of incorrectly labeled samples outweighs ignoring difficult ones.
@@
We instead determine these samples by first running the system with the base Llama 3 model, with a prompt indicating to always produce a prediction for each of the 4 candidate retrieval approaches mentioned previously. We use GPT-4o to evaluate whether any of the generated answers are correct. If any are correct, the original label is retained for the training, otherwise ”i don’t know” is used as the target label. In the case of false premise questions, we always retain the original label as the target label. We repeat this process for each task, given that each has access to a different data sources, to generate a training dataset for each LoRa adapter.
@@
We demonstrate the effectiveness of relabeling in Table 2. We ran 3 different answer generation setups for the 500 sample Task 1 hold out set we created. The first is an unmodified Llama 3 8B model, the second is a LoRa model using the original targets, and the final a LoRa model with relabeled targets. As shown, using the original targets provide the best accuracy, but also worsens hallucinations over the base model. While using the relabeled targets hurts accuracy, it also substantially reduces hallucinations, providing the best CRAG Score amongst the three.
@@ 5. Results
The performance difference between head, torso, and tail questions appeared less substantial than our original expectations, though clearly performance drops off as popularity falls. Interestingly, Task 3 here under performs the other tasks in head, torso, and tail. We suspect that including substantially more search results includes overlapping entities / subjects, at which point conflicting information would be difficult to resolve.
@@ 6. Future Work
Additionally, we hope to explore larger Llama models, 70B+, in the future for this task. We were unable to get the 70B model running in the competition compute environments, so did not spend much timer looking at larger models. However, it is very likely moving to larger models would provide a substantial improvement over the 8B model.
@@ 7. Conclusion
In this work we presented MARAGS, a multi-adapter solution to the CRAG dataset. We demonstrated the effectiveness of training individual LoRa adapters for the 4 tasks in the pipeline, specifically API call generation and Task 1,2, and 3 answer generation. CRAG presents a variety of different tasks and questions to allow the tracking of progress of various methods used to build RAG systems. The penalization of hallucinations is a unique and important feature as future AI systems become increasingly common throughout society, as hallucinations hurt user trust in these systems. We discussed our methods for reducing these hallucinations, but they are not without cost, as in some cases the model fails to output previously known knowledge. Clearly the importance of balancing these two factors is a key to leveraging LLMs to their full potential, while also improving user trust.

### v2

@@ 1. Introduction
Retrieval augmented generation (RAG) has been a popular approach for question answering systems for some time (Lewis et al., 2020), although recently has become a very popular approach for a wide range of tasks due to the zero-shot capabilities of large language models (LLMs) with an appropriate prompt and access to the relevant context for the task. Despite the existence of numerous question answering benchmarks, many do accurately reflect the diverse usage of current RAG systems. Thus, tracking both the efficacy of certain RAG architectures as well as tracking process remains difficult. The CRAG (Yang et al., 2024) benchmark aims to resolve this with 3 different subtasks representing realistic RAG usage scenarios. The final key element of the CRAG benchmark is its scoring metric which explicitly punishes hallucinations. With the rising capabilities of LLMs, increasingly their outputs are taken at face value, despite the known issue of hallucinations. This has led to high profile incidents causing concern with their use (law, [n. d.]). The CRAG score aims to punish hallucinated answers and encourages returning missing answers, equivalent to returning ”i don’t know” from the model, by giving scores of 1, 0, and -1 to correct, missing, and hallucinated answers respectively. To address these various tasks, we train individual adapters for each of the various tasks, as well as API call generation required for accessing information in the mock API. This approach allows us to use a single LLama 3 (Dubey et al., 2024) in memory while swapping out adapters based on the current needs.
@@ 2. Related Works
The initial approach of Lewis et al. (Lewis et al., 2020) showed the benefits of providing additional text context for seq2seq models for NLP tasks that are knowledge intensive. Using BART, they were able to improve question answering tasks using dual biencoders for retrieval and training the model jointly, without the need for knowing which documents were relevant.
Adapters have become increasingly used since introduced by Houlsby et al. (Houlsby et al., 2019). LoRa (Hu et al., 2022) has become a popular adapter approach, particularly for LLMs as they have grown substantially larger in recent years. The use of adapters allows modifying a model’s output without training the entire network, which substantially saves on VRAM memory when training. Hu et al. (Hu et al., 2022) discovered that when replacing a dot product between large vectors with an intermediate dot product of a much lower rank vector, the impacts on performance were minimal while further reducing the training parameters required.
@@ 3. CRAG Dataset
CRAG is a question answering dataset aimed at providing a realistic task to benchmark RAG systems as they are used in practice with a diverse set of questions, including 8 distinct question types and 5 distinct domains. Additionally, two sources of diversity which pose difficulty for LLMs are how dynamic a question’s answer is and the popularity of the topic of the question. As shown in the baseline systems, real-time answers pose a challenge for RAG systems and they similarly struggle when the topic of the question is less common (referred to as ”torso” and ”tail” questions).
@@ 3.3. Task 3
Finally, Task 3 represents an extension of Task 2, where the system has access to both HTML and the knowledge graph API. However, in this task, the number of HTML documents to be processed is 50. This task is meant to measure both the computational efficiency of the approach as well as its ability to filter large amounts of potentially irrelevant information.
@@ 4.1. Webpage Processing
In our case, we utilize the structure of the HTML to provide the segmentation. We traverse the tree structure of the parse HTML with breadth-first search. Any time the length of the text contained within a node (which includes all of the text of its descendants) is less than 2000 characters, we treat that as a segment. If a leaf node is reached with greater than 2000 characters, the text is split on the white space and into as many segments are needed such that each segment is under the threshold. The segment length was determined via inspection of HTML documents and their associated segmentation, thus future work could treat this as a hyperparameter and tune for performance.
@@ 4.2. API Call Generation
We acknowledge that this approach to annotation is not optimal, as it likely results in successful, but incorrectly selected API endpoint calls. However, manually annotating each question to determine the correct API call and validating the returned information were indeed relevant would have been too time consuming given the size of the dataset.
@@ 4.3. Candidate Ranking
For candidate ranking, we attempted 4 different candidate ranking approaches. We utilized TF-IDF, a biencoder, cross-encoder, and an ensemble of the mentioned approaches (using mean rank as the ranking metric). Our TF-IDF implementation is based on Scikit-Learn (Pedregosa et al., 2011). The biencoder and cross-encoder are from the SentenceTransformer (Reimers and Gurevych, 2019) library, specifically the ”multi-qa-MiniLM-L6-cos-v1” 111https://huggingface.co/sentence-transformers/multi-qa-MiniLM-L6-cos-v1 and ”ms-marco-MiniLM-L-6-v2” 222https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2 respectively.
Evaluating candidate ranking in isolation is difficult, as relevant information is not labeled, so using the system accuracy and CRAG score is the most straight forward way to compare differences in each system. However, to test the various systems, we use the base Llama-3 8B model with no adapters for each retrieval approach and use the accuracy metric to determine the best performing approach. We use accuracy instead of CRAG at this stage for ranking, as we think this is a better representation of how often the relevant information retrieved. For a test set, we randomly select 500 samples from the Task 1 dataset.
@@
The results of this experiment are shown in Table 1. From the results, the cross-encoder is the best performing system, thus we used it for our retriever. We suspect that with proper tuning TF-IDF and ensembling would be much more performant overall, but as mentioned running extensive experimentation is difficult as it requires LLM generation to get an overall accuracy score. Using an LLM to label passages as relevant or not is a possible approach to allow for tuning of just the retriever, however we did not explore this.
Despite the cross-encoder being the most computationally expensive approach, we found it to be fast enough for processing the candidates in the required 30 seconds per sample. In the case of Task 3, it was necessary to use Python’s multiprocessing library to process multiple HTMLs simultaneously to meet the runtime requirement.
@@ 4.4. Retrieval Augmented Generation
Our initial prompt structure started with the question first, then all of the retrieved candidates prior to the Llama model response, however we noticed that often due to how much context would be provided, the model would occasionally forget the question being asked. For example a question like ”What is the elevation of Honolulu?” would result in an answer of ”Honolulu population is 343,421”, indicating the model remembered the subject of the question, but incorrectly answered with the population, rather than elevation. Our subsequent prompt structure placed the question after the context, which resolved the issue.
For training Llama 3, we trained LoRa models for each task individually. Given the penalization of hallucinations in the scoring metric, we try to take steps to mitigate further hallucinations due to fine-tuning, as it has been observed that fine-tuning LLMs can be a source of further hallucinations (Gekhman et al., 2024). This likely applies to RAG systems in cases where the expected answer is not answerable given the context provided, i.e. no relevant information is given and the question is not answerable without further context, yet the model is trained to output the target answer regardless. Thus for our training setup, we first relabel the target for training samples in cases where our candidate retrieval system likely does not provide the correct information and Llama does not know the answer without that information. We use the provided dev set for training, with the 500 set sample used for retrieval comparison treated as our holdout set.
@@
| - LoRa | 0.398 | 0.602 | -0.204 |
| - LoRa (relabeled) | 0.242 | 0.056 | 0.186 |
Our initial approach for determining which training samples need relabeling has been explored previously (Wang et al., 2018a). A common and simple approach to filter/relabel incorrectly333”Incorrectly” here simply means in the context of our retrieval system, not that the provided answer is not true. labeled samples is to use a particular sample’s training loss after training to the point of over-fitting. High loss examples after over-fitting likely indicate examples that are incorrectly labeled and thus can be filtered out. Not all samples with high loss will be incorrect labels, instead simply being hard examples, yet typically the benefit of disposing of incorrectly labeled samples outweighs ignoring difficult ones.
@@
We instead determine these samples by first running the system with the base Llama 3 model, with a prompt indicating to always produce a prediction for each of the 4 candidate retrieval approaches mentioned previously. We use GPT-4o to evaluate whether any of the generated answers are correct. If any are correct, the original label is retained for the training, otherwise ”i don’t know” is used as the target label. In the case of false premise questions, we always retain the original label as the target label. We repeat this process for each task, given that each has access to different data sources, to generate a training dataset for each LoRa adapter.
@@
We demonstrate the effectiveness of relabeling in Table 2. We ran 3 different answer generation setups for the 500 sample Task 1 hold out set we created. The first is an unmodified Llama 3 8B model, the second is a LoRa model using the original targets, and the final a LoRa model with relabeled targets. As shown, using the original targets provides the best accuracy, but also worsens hallucinations over the base model. While using the relabeled targets hurts accuracy, it also substantially reduces hallucinations, providing the best CRAG Score among the three.
@@ 5. Results
The performance difference between head, torso, and tail questions appeared less substantial than our original expectations, though clearly performance drops off as popularity falls. Interestingly, Task 3 underperforms the other tasks in head, torso, and tail. We suspect that including substantially more search results includes overlapping entities / subjects, at which point conflicting information would be difficult to resolve.
@@ 6. Future Work
Additionally, we hope to explore larger Llama models, 70B+, in the future for this task. We were unable to get the 70B model running in the competition compute environments, so did not spend much time looking at larger models. However, it is very likely moving to larger models would provide a substantial improvement over the 8B model.
@@ 7. Conclusion
In this work we presented MARAGS, a multi-adapter solution to the CRAG dataset. We demonstrated the effectiveness of training individual LoRa adapters for the 4 tasks in the pipeline, specifically API call generation and Task 1,2, and 3 answer generation. CRAG presents a variety of different tasks and questions to allow the tracking the progress of various methods used to build RAG systems. The penalization of hallucinations is a unique and important feature as future AI systems become increasingly common throughout society, as hallucinations hurt user trust in these systems. We discussed our methods for reducing these hallucinations, but they are not without cost, as in some cases the model fails to output previously known knowledge. Clearly the importance of balancing these two factors is a key to leveraging LLMs to their full potential, while also improving user trust.